# A Multi-fidelity Estimator of the Expected Information Gain for Bayesian Optimal Experimental Design

### v1

@@ 4.2 The MF-EIG estimator
$u_{m}(\xi,z^{(i)})\approx u_{\mathcal{E}}(\xi,z^{(i)}),\quad m=0,\ldots,M.$ (23)
@@
$\tilde{U}(\xi,z;\alpha,\mathcal{A}):=\hat{U}_{0}\left(\xi,z_{0}\right)+\sum^{M}_{m=1}\alpha_{m}\left(\hat{U}_{m}\left(\xi,z_{m}^{\ast}\right)-\hat{U}_{m}\left(\xi,z_{m}\right)\right),$ (24)
@@ 4.3 MF-EIG hyperparameters
$C=\mathbb{C}\text{ov}\left[\mathrm{u}(\xi,Z)\right]\approx\hat{\Sigma}(\xi,z_{\text{pilot}}):=\frac{1}{N_{\text{pilot}}-1}\sum_{j=1}^{N_{\text{pilot}}}\left(\mathrm{u}(\xi,z^{(j)})-\overline{\mathrm{u}}(\xi,z_{\text{pilot}})\right)\left(\mathrm{u}(\xi,z^{(j)})-\overline{\mathrm{u}}(\xi,z_{\text{pilot}})\right)^{\intercal},$ (26)
@@
$\overline{\Sigma}:=\frac{1}{N_{\xi}}\sum_{k=1}^{N_{\xi}}\hat{\Sigma}(\xi_{k}),$ (27)
where $\xi_{k}$ are $N_{\xi}$ designs selected for averaging (e.g., sampled uniformly in $\Xi$). Additionally, the pilot samples can be used to estimate the average computational costs (e.g., CPU-hours of run time) of each utility model, if not already known. Once the covariance matrix and model costs are determined, the optimization problem in (16) and (17) is solved. The resulting hyperparameter values are then applied to all MF-EIG estimators across the design domain.
@@ 4.4 NMC-style utility models
We now provide a detailed presentation of MF-EIG estimators based on NMC-style utility models. Similar to before, subscript 00 denotes the high-fidelity target quantities introduced in Sections 2 and 4.1; that is, $Y_{0}=Y$, $h_{0}=h$, $J_{0}=J$, $p_{Y_{0}|\theta,\xi}(y_{0}|\theta,\xi)=p_{Y|\theta,\xi}(y|\theta,\xi)$, $p_{Y_{0}|\xi}(y_{0}|\xi)=p_{Y|\xi}(y|\xi)$, and $u_{\mathcal{E},0}=u_{\mathcal{E}}$. Next, for $m=1,\ldots,M$, we introduce low-fidelity injective and continuously invertible data models $h_{m}$ relating the common random data noise $\mathcal{E}$ to low-fidelity observation variables $Y_{m}$, where $Y_{m}=h_{m}(\mathcal{E};\Theta,\xi)$ with inverse $\mathcal{E}=h_{m}^{-1}(Y_{m};\Theta,\xi)$ and Jacobian $J_{m}=\frac{\partial h_{m}}{\partial\epsilon}$ with inverse $J_{m}^{-1}=\frac{\partial h_{m}^{-1}}{\partial y_{m}}$. Each $Y_{m}$ is also associated with its own low-fidelity likelihood and evidence, $p(y_{m}|\theta,\xi)$ and $p(y_{m}|\xi)$. Similar to (18) and (19), they can be transformed into
$p(y_{m}|\theta,\xi)$ $=p_{\mathcal{E}|\xi}\left(h_{m}^{-1}(y_{m};\theta,\xi)\big{|}\xi\right)\,\big{\lvert}J_{m}^{-1}(y_{m};\theta,\xi)\big{\rvert}=p(\epsilon|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert},$ (28)
$p(y_{m}|\xi)$ $=\int_{\widetilde{\Theta}}p(y_{m}|\tilde{\theta},\xi)\,p(\tilde{\theta})\,\text{d}{\tilde{\theta}}=\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}|\xi)\,\big{\lvert}J_{m}^{-1}(y_{m};\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta},$ (29)
@@
$u_{\mathcal{E},m}(\xi,\epsilon^{(i)},\theta^{(i)}):=\log\left[\frac{p(\epsilon^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon^{(i)};\theta^{(i)},\xi);\theta^{(i)},\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}({\epsilon}^{(i)};{\theta}^{(i)},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right].$ (30)
@@
$u_{m}(\xi,z^{(i)}):=\log{\left[\frac{p(\epsilon^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon^{(i)};\theta^{(i)},\xi);\theta^{(i)},\xi)\big{\rvert}}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}({\epsilon}^{(i)};\theta^{(i)},\xi);\tilde{\theta}^{(i,j)},\xi)\big{\rvert}}\right]},\quad m=0,\ldots,M,$ (31)
where $\tilde{\theta}^{(i,j)}$ are $N_{\text{in},m}$ i.i.d. samples from the prior $p(\theta)$. In summary, $u_{m}$ for $m=0,\ldots,M$ are respective approximations to $u_{\mathcal{E},m}$ resulting from the inner-loop MC, and $u_{\mathcal{E},m}$ for $m=1,\ldots,M$ are all approximations to the exact utility $u_{\mathcal{E},0}=u_{\mathcal{E}}$ due to their low-fidelity data models $h_{m}$.
@@ 4.4.1 Additive independent noise
$u_{m}(\xi,z^{(i)})$ $=\log{\left[\frac{p(\epsilon^{(i)}|\xi)}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)}\right]}=\log{\left[\frac{p(\epsilon^{(i)}|\xi)}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}\left(g_{m}(\theta^{(i)},\xi)+\epsilon^{(i)}-g_{m}(\tilde{\theta}^{(i,j)},\xi)\big{|}\xi\right)}\right]}$ (32)
@@ 4.4.2 Scaled noise
$=\log\left[\frac{{p(\epsilon^{(i)})}\,\Big{/}\,\left\lvert\text{diag}(g_{m}(\theta^{(i)},\xi))\right\rvert}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}\left((g_{m}(\theta^{(i)},\xi)+g_{m}(\theta^{(i)},\xi)\circ\epsilon^{(i)}-g(\tilde{\theta}^{(i,j)},\xi))\oslash g(\tilde{\theta}^{(i,j)},\xi)\right)\,\Big{/}\,\left\lvert\text{diag}(g_{m}(\tilde{\theta}^{(i,j)},\xi))\right\rvert}\right]$ (33)
@@ 4.5.2 Inner-loop sample sizes
$\min_{\mathrm{N}_{\text{in}}\in\mathbb{N}^{M}}\quad\left\{\quad\begin{aligned} \min_{\mathcal{A}\in\mathbb{A}}&\quad\mathbb{V}\text{ar}[\tilde{U}(\xi,Z;\alpha^{\ast},\mathcal{A},\mathrm{N}_{\text{in}})]\\ \text{subject to}&\quad\mathcal{W}(w{(\mathrm{N}_{\text{in}}}),\mathcal{A})\leq w_{\text{budget}}\end{aligned}\quad\right\},$ (34)
@@ 5.1 Case 1: nonlinear benchmark
$Y_{0}=g_{0}(\Theta,\xi)+\mathcal{E}=\Theta^{3}\xi^{2}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$ (35)
@@
$Y_{1}$ $=g_{1}(\Theta,\xi)+\mathcal{E}=k_{1}\,\Theta^{2.5}\xi^{1.75}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$ (36)
$Y_{2}$ $=g_{2}(\Theta,\xi)+\mathcal{E}=k_{2}\,\Theta^{2}\xi^{1.5}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$ (37)
@@ 5.1.1 Inner-loop sample sizes
We consider two approaches for determining the inner-loop sample sizes (no sample reuse). In the first approach, the inner-loop sample sizes are naïvely chosen to match that of the high-fidelity model, i.e., $N_{\text{in},0}=N_{\text{in},1}=N_{\text{in},2}=2500$. Any improvement in the estimator’s performance thus arises solely from the multi-fidelity extension to the single-fidelity NMC. The optimal allocation $\mathcal{A}$ is found to be a generalized version of ACVIS (independent samples) [22] called the GISSR estimator [8]. Table 1 presents the MF-EIG estimator variance and variance reduction ratio relative to the single-fidelity NMC estimator, averaged over 41 evenly-spaced designs across $\Xi$. Projected values are predicted by MXMCPy during hyperparameter optimization, and empirical values are obtained from repeating 50 estimation trials using fresh samples. As shown by the first two columns of Table 1, even with this naïve-$\text{N}_{\text{in}}$ approach, 6–8 times variance reduction is achieved relative to single-fidelity NMC.
@@
Variance (projected) $6.34\times 10^{-4}$ $9.66\times 10^{-5}$ $8.61\times 10^{-5}$ $3.70\times 10^{-5}$ $3.37\times 10^{-5}$
@@
In the second approach, the inner-loop sample sizes are optimized following (34). The inner optimization, under a given $\mathrm{N}_{\text{in}}$, is solved using MXMCPy. The outer optimization, an integer program, is tackled using a grid search over reasonable values of $\mathrm{N}_{\text{in}}$; here we use a $50\times 50$ grid over an integer-valued search space of $[25,4000]^{2}$. Grid search, however, is only practical when $M$ is relatively small and where the inner optimization is relatively fast. Advanced techniques of integer programming, and mixed-integer strategies that combine the inner and outer optimization problems, can be quite useful but are not explored in this paper. The optimal allocation $\mathcal{A}$ is found to be a generalized version of MLMC [19] called the WRDIFF estimator [8].
@@
(a) Comp. cost (dashed) and correlation to $u_{0}$ (solid)
@@
Figure 1: Case 1 (no reuse). Computational cost and design-averaged correlation as $N_{\text{in},1}$ and $N_{\text{in},2}$ change.
Figure 2: Case 1 (no reuse). Projected MF-EIG estimator variance across $\mathrm{N}_{\text{in}}$. The contour levels use a power law normalization for enhanced visualization.
Figure 3 shows the empirical estimator mean and ±$\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and the optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators. In obtaining these results, we employ the common random number technique by fixing the random seed across $\Xi$ when sampling $Z$ in the ACV outer loop—that is, the $z$ input on the left-hand side of (24) is the same for every $\xi$ queried. This in turn leads to increased correlation among estimators across $\Xi$ and in turn improving the smoothness of these curves. Overall, the estimator standard deviation is much smaller for MF-EIG compared to single-fidelity NMC. Notably, the estimator standard deviation increases at larger values of $\xi$, suggesting that MF-EIG may be further improved from $\xi$-specific tuning. The excellent alignment in the mean curves confirms the unbiasedness of the MF-EIG estimator with respect to the high-fidelity mean. The design-averaged optimal-$\text{N}_{\text{in}}$ estimator variance and variance reduction ratio relative to single-fidelity NMC, both projected and empirical, are summarized in the third column of Table 1. Optimal-$\text{N}_{\text{in}}$ achieves around 7–9 times variance reduction relative to single-fidelity NMC, and a 6–12% design-averaged improvement over naïve-$\text{N}_{\text{in}}$.
Figure 3: Case 1 (no reuse). Empirical estimator mean and ±$\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators.

### v2

@@ 4.2 The MF-EIG estimator
(23) $u_{m}(\xi,z^{(i)})\approx u_{\mathcal{E}}(\xi,z^{(i)}),\quad m=0,\ldots,M.$
@@
(24) $\tilde{U}(\xi,z;\alpha,\mathcal{A}):=\hat{U}_{0}\left(\xi,z_{0}\right)+\sum^{M}_{m=1}\alpha_{m}\left(\hat{U}_{m}\left(\xi,z_{m}^{\ast}\right)-\hat{U}_{m}\left(\xi,z_{m}\right)\right),$
@@ 4.3 MF-EIG hyperparameters
(26) $C$ $=\mathbb{C}\text{ov}\left[\mathrm{u}(\xi,Z)\right]$
$\approx\hat{\Sigma}(\xi,z_{\text{pilot}}):=\frac{1}{N_{\text{pilot}}-1}\sum_{j=1}^{N_{\text{pilot}}}\left(\mathrm{u}(\xi,z^{(j)})-\overline{\mathrm{u}}(\xi,z_{\text{pilot}})\right)\left(\mathrm{u}(\xi,z^{(j)})-\overline{\mathrm{u}}(\xi,z_{\text{pilot}})\right)^{\intercal},$
@@
(27) $\overline{\Sigma}:=\frac{1}{N_{\xi}}\sum_{k=1}^{N_{\xi}}\hat{\Sigma}(\xi_{k}),$
where $\xi_{k}$ are $N_{\xi}$ designs selected for averaging (e.g., sampled uniformly in $\Xi$). Additionally, the pilot samples can be used to estimate the average computational costs (e.g., CPU-hours of run time) of each utility model, if not already known. Once the covariance matrix and model costs are determined, the optimization problem in (16) and (17) is solved. The resulting hyperparameter values are then applied to all MF-EIG estimators across the design domain. Adopting a single MF-EIG estimator structure and using uniform pilot sample sizes across the entire design domain is practical but may be suboptimal when the covariance matrix varies significantly throughout the domain. A more tailored approach could involved redesigning the MF-EIG estimator and allocating more or fewer pilot samples at different design locations, potentially increasing the total number of pilot samples needed to accurately estimate local covariance matrix. Alternatively, applying smoothing or surrogate modeling techniques to the covariance matrix across the design space may reduce the need for additional pilot samples. We leave the exploration of these strategies to future work.
@@ 4.4 NMC-style utility models
We now provide a detailed presentation of MF-EIG estimators based on NMC-style utility models. Similar to before, subscript 0 denotes the high-fidelity target quantities introduced in Sections 2 and 4.1; that is, $Y_{0}=Y$, $h_{0}=h$, $J_{0}=J$, $p_{Y_{0}|\theta,\xi}(y_{0}|\theta,\xi)=p_{Y|\theta,\xi}(y|\theta,\xi)$, $p_{Y_{0}|\xi}(y_{0}|\xi)=p_{Y|\xi}(y|\xi)$, and $u_{\mathcal{E},0}=u_{\mathcal{E}}$. Next, for $m=1,\ldots,M$, we introduce low-fidelity injective and continuously invertible data models $h_{m}$ relating the common random data noise $\mathcal{E}$ to low-fidelity observation variables $Y_{m}$, where $Y_{m}=h_{m}(\mathcal{E};\Theta,\xi)$ with inverse $\mathcal{E}=h_{m}^{-1}(Y_{m};\Theta,\xi)$ and Jacobian $J_{m}=\frac{\partial h_{m}}{\partial\epsilon}$ with inverse $J_{m}^{-1}=\frac{\partial h_{m}^{-1}}{\partial y_{m}}$. Each $Y_{m}$ is also associated with its own low-fidelity likelihood and evidence, $p(y_{m}|\theta,\xi)$ and $p(y_{m}|\xi)$. Similar to (18) and (19), they can be transformed into
(28) $p(y_{m}|\theta,\xi)$ $=p_{\mathcal{E}|\xi}\left(h_{m}^{-1}(y_{m};\theta,\xi)\big{|}\xi\right)\,\big{\lvert}J_{m}^{-1}(y_{m};\theta,\xi)\big{\rvert}=p(\epsilon|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon;\theta,\xi);\theta,\xi)\big{\rvert}$,
(29) $p(y_{m}|\xi)$ $=\int_{\widetilde{\Theta}}p(y_{m}|\tilde{\theta},\xi)\,p(\tilde{\theta})\,\text{d}{\tilde{\theta}}=\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}|\xi)\,\big{\lvert}J_{m}^{-1}(y_{m};\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}$,
@@
(30) $u_{\mathcal{E},m}(\xi,\epsilon^{(i)},\theta^{(i)}):=\log\left[\frac{p(\epsilon^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon^{(i)};\theta^{(i)},\xi);\theta^{(i)},\xi)\big{\rvert}}{\int_{\widetilde{\Theta}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}({\epsilon}^{(i)};{\theta}^{(i)},\xi);\tilde{\theta},\xi)\big{\rvert}\,p(\tilde{\theta})\,\text{d}\tilde{\theta}}\right$].
@@
(31) $u_{m}(\xi,z^{(i)}):=\log{\left[\frac{p(\epsilon^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}(\epsilon^{(i)};\theta^{(i)},\xi);\theta^{(i)},\xi)\big{\rvert}}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)\,\big{\lvert}J_{m}^{-1}(h_{m}({\epsilon}^{(i)};\theta^{(i)},\xi);\tilde{\theta}^{(i,j)},\xi)\big{\rvert}}\right]}$,
for $m=0,\ldots,M$, and where $\tilde{\theta}^{(i,j)}$ are $N_{\text{in},m}$ i.i.d. samples from the prior $p(\theta)$. In summary, $u_{m}$ for $m=0,\ldots,M$ are respective approximations to $u_{\mathcal{E},m}$ resulting from the inner-loop MC, and $u_{\mathcal{E},m}$ for $m=1,\ldots,M$ are all approximations to the exact utility $u_{\mathcal{E},0}=u_{\mathcal{E}}$ due to their low-fidelity data models $h_{m}$.
@@ 4.4.1 Additive independent noise
(32) $u_{m}(\xi,z^{(i)})$ $=\log{\left[\frac{p(\epsilon^{(i)}|\xi)}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}(\tilde{\epsilon}_{m}^{(i)}|\xi)}\right]}$
$=\log{\left[\frac{p(\epsilon^{(i)}|\xi)}{\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}p_{\mathcal{E}|\xi}\left(g_{m}(\theta^{(i)},\xi)+\epsilon^{(i)}-g_{m}(\tilde{\theta}^{(i,j)},\xi)\big{|}\xi\right)}\right]}$
@@ 4.4.2 Scaled noise
(33) $=\log\left[\frac{\frac{\displaystyle{p(\epsilon^{(i)}|\xi)}}{\displaystyle\left\lvert\text{diag}(g_{m}(\theta^{(i)},\xi))\right\rvert}}{\displaystyle\frac{1}{N_{\text{in},m}}\sum_{j=1}^{N_{\text{in},m}}\frac{p_{\mathcal{E}|\xi}\left((g_{m}(\theta^{(i)},\xi)+g_{m}(\theta^{(i)},\xi)\circ\epsilon^{(i)}-g(\tilde{\theta}^{(i,j)},\xi))\oslash g(\tilde{\theta}^{(i,j)},\xi)\Big{|}\xi\right)}{\displaystyle\left\lvert\text{diag}(g_{m}(\tilde{\theta}^{(i,j)},\xi))\right\rvert}}\right$]
@@ 4.5.2 Inner-loop sample sizes
(34) $\min_{\mathrm{N}_{\text{in}}\in\mathbb{N}^{M}}\quad\left\{\quad\begin{aligned} \min_{\mathcal{A}\in\mathbb{A}}&\quad\mathbb{V}\text{ar}[\tilde{U}(\xi,Z;\alpha^{\ast},\mathcal{A},\mathrm{N}_{\text{in}})]\\ \text{subject to}&\quad\mathcal{W}(w{(\mathrm{N}_{\text{in}}}),\mathcal{A})\leq w_{\text{budget}}\end{aligned}\quad\right\},$
@@ 5.1 Case 1: nonlinear benchmark
(35) $Y_{0}=g_{0}(\Theta,\xi)+\mathcal{E}=\Theta^{3}\xi^{2}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$
@@
(36) $Y_{1}$ $=g_{1}(\Theta,\xi)+\mathcal{E}=k_{1}\,\Theta^{2.5}\xi^{1.75}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$
(37) $Y_{2}$ $=g_{2}(\Theta,\xi)+\mathcal{E}=k_{2}\,\Theta^{2}\xi^{1.5}+\Theta\exp{(-|0.2-\xi|)}+\mathcal{E},$
@@ 5.1.1 Inner-loop sample sizes
We consider two approaches for determining the inner-loop sample sizes (no sample reuse). In the first approach, the inner-loop sample sizes are naïvely chosen to match that of the high-fidelity model, i.e., $N_{\text{in},0}=N_{\text{in},1}=N_{\text{in},2}=2500$. Any improvement in the estimator’s performance thus arises solely from the multi-fidelity extension to the single-fidelity NMC. The optimal allocation $\mathcal{A}$ is found to be a generalized version of ACVIS (independent samples) [22] called the GISSR estimator [8]. Table 1 presents the MF-EIG estimator variance and variance reduction ratio relative to the single-fidelity NMC estimator, averaged over 41 evenly-spaced designs across $\Xi$. The MXMCPy tool takes just 1.38 seconds to find the optimal MF-EIG hyperparameters on a single 3.0 GHz Intel Xeon Gold 6154 processor. Projected values are predicted by MXMCPy during hyperparameter optimization, and empirical values are obtained from repeating 50 estimation trials using fresh samples. As shown by the first two columns of Table 1, even with this naïve-$\text{N}_{\text{in}}$ approach, 6–8 times variance reduction is achieved relative to single-fidelity NMC.
@@
Variance (projected) $6.34\times 10^{-4}$ $9.66\times 10^{-5}$ $8.61\times 10^{-5}$ $70\times 10^{-5} 3.3$$37\times 10^{-5}$
@@
In the second approach, the inner-loop sample sizes are optimized following (34). The inner optimization, under a given $\mathrm{N}_{\text{in}}$, is solved using MXMCPy. The outer optimization, an integer program, is tackled using a grid search over reasonable values of $\mathrm{N}_{\text{in}}$; here we use a $50\times 50$ grid over an integer-valued search space of $25,4000]^{2}$. Grid search, however, is only practical when $M$ is relatively small and where the inner optimization is relatively fast. Advanced techniques of integer programming, and mixed-integer strategies that combine the inner and outer optimization problems, can be quite useful but are not explored in this paper. The optimal allocation $\mathcal{A}$ is found to be a generalized version of MLMC [19] called the WRDIFF estimator [8].
@@
(a) Computational cost (dashed) and correlation to $u_{0}$ (solid)
@@
Figure 1: Case 1 (no reuse). Computational cost and design-averaged correlation as $N_{\text{in},1}$ and $N_{\text{in},2}$ change. (a) Low-fidelity log-ratio model computational costs (dashed lines) and correlations (solid lines) to the high-fidelity log-ratio model over inner-loop sample sizes (no sample reuse in the inner loop). The complex trade-off between correlations and costs dictate the best inner-loop sample sizes to use. (b) Correlations between low-fidelity log-ratio model over inner-loop sample sizes (no sample reuse in the inner loop).
Figure 2: Case 1 (no reuse). Projected MF-EIG estimator variance across different possible low-fidelity inner-loop sample sizes,
$\mathrm{N}_{\text{in}}$. The contour levels use a power law normalization for enhanced visualization.
Figure 3 shows the empirical estimator mean and $\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and the optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators. In obtaining these results, we employ the common random number technique by fixing the random seed across $\Xi$ when sampling $Z$ in the ACV outer loop—that is, the $z$ input on the left-hand side of (24) is the same for every $\xi$ queried. This in turn leads to increased correlation among estimators across $\Xi$ and in turn improving the smoothness of these curves. Overall, the estimator standard deviation is much smaller for MF-EIG compared to single-fidelity NMC. Notably, the estimator standard deviation increases at larger values of $\xi$, suggesting that MF-EIG may be further improved from $\xi$-specific tuning. The excellent alignment in the mean curves confirms the unbiasedness of the MF-EIG estimator with respect to the high-fidelity mean. The design-averaged optimal-$\text{N}_{\text{in}}$ estimator variance and variance reduction ratio relative to single-fidelity NMC, both projected and empirical, are summarized in the third column of Table 1. Optimal-$\text{N}_{\text{in}}$ achieves around 7–9 times variance reduction relative to single-fidelity NMC, and a 6–12% design-averaged improvement over naïve-$\text{N}_{\text{in}}$.
Figure 3: Case 1 (no reuse). Empirical estimator mean and $\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators.